OpenStudy (anonymous):

find the domain of f of f(x)=x/(sqrtx^2-4) PLEASE HELP!

OpenStudy (alfie):

\[f(x)=\frac{x}{\sqrt x^2-4}\] Like this? Or: \[f(x)=\frac{x}{\sqrt {x^2 -4}}\] I guess the second one, so, denominator different from zero, everything under the root greater or equal to zero: \[\sqrt {x^2-4} \neq 0\] \[x^2-4 \ge 0\] Notice, you can merge the two saying x^2-4 > 0, but you'll find it out solving :)

jhonyy9 (jhonyy9):

x can have values from R without -2 and 2